OpenStudy (anonymous):

find the points of inflection and identify where the graph of f(x)=x+2cosx is concave up or concave down on the interval [0,2pi]. find the limit approaching positive infinity of x-cosx/x

OpenStudy (anonymous):

\[f(x)=x+2\cos x~~\Rightarrow~~f'(x)=1-2\sin x~~\Rightarrow~~f''(x)=-2\cos x\] Solve for \(x\) with \(0\le x\le2\pi\): \[-2\cos x=0\] You should get two solutions. For the limit: \[\lim_{x\to\infty}\frac{x-\cos x}{x}=\frac{\infty}{\infty}\] Apply L'Hopital's rule: \[\lim_{x\to\infty}\frac{1-\sin x}{1}=\frac{1-(\pm1)}{1}\] The oscillating nature says the limit doesn't exist.